### v1

# REACT3D: Recovering Articulations for Interactive Physical 3D Scenes

Abstract
Interactive 3D scenes are increasingly vital for embodied intelligence, yet existing datasets remain limited due to the labor-intensive process of annotating part segmentation, kinematic types, and motion trajectories. We present REACT3D, a scalable zero-shot framework that converts static 3D scenes into simulation-ready interactive replicas with consistent geometry, enabling direct use in diverse downstream tasks. Our contributions include: (i) openable-object detection and segmentation to extract candidate movable parts from static scenes, (ii) articulation estimation that infers joint types and motion parameters, (iii) hidden-geometry completion followed by interactive object assembly, and (iv) interactive scene integration in widely supported formats to ensure compatibility with standard simulation platforms. We achieve state-of-the-art performance on detection/segmentation and articulation metrics across diverse indoor scenes, demonstrating the effectiveness of our framework and providing a practical foundation for scalable interactive scene generation, thereby lowering the barrier to large-scale research on articulated scene understanding. Our project page is react3d.github.io.
I INTRODUCTION
High-fidelity, interactive 3D assets are critical for a wide range of applications. They enable immersive experiences in virtual and mixed reality, support realistic content creation for gaming and film production, and facilitate the development of autonomous robotic systems by serving as training environments for navigation and manipulation tasks [1]. These applications demand large-scale 3D datasets that offer both photorealistic rendering and physically plausible interactions, such as object picking, placing, or articulating structures like doors and drawers. Automating the generation of such datasets is essential to scale up these efforts effectively.
Significant effort has been dedicated to the automatic generation of 3D assets. Early works primarily focused on the creation of static 3D assets, often converting real-world scans or image sequences into accurate 3D representations such as meshes or point clouds [2]. These efforts have led to the release of several high-quality static 3D datasets. More recently, the rapid advancement of generative models has enabled the synthesis of large-scale static scenes with diverse modalities and formats [3]. As a result, the generation of static 3D scenes, whether through real-world reconstruction or generative modeling, has become increasingly mature.
Figure 1: REACT3D transforms static 3D scenes into interactive scenes in a zero-shot manner. The generated interactive scenes are spatially aligned with the static input and preserve the original geometry and appearance. Our results are readily compatible with multiple simulation platforms, supporting diverse downstream tasks such as robotic perception, interaction, and embodied intelligence.
In parallel, the research community has begun to explore the generation of dynamic scenes, either by reconstructing real-world interaction sequences [4] or by manually designing 3D environments with humans in the loop. However, despite promising progress, the scalability and quality of interactive scene generation still lag behind those of static scene generation.
In this work, we are motivated by the growing need for large-scale, interactive 3D environments to support embodied AI research. We propose a novel, application-driven framework for converting high-quality static scenes into physics-aware, interactive environments (Fig. 1). Our system leverages recent advances in vision foundation models to enable zero-shot transfer from standard 3D scene formats (e.g., .ply, .glb) to simulation-ready formats such as URDF and USD. These converted scenes support physical simulation and enable robotic agents to navigate and interact with articulated objects. Our method provides a reliable and efficient solution for obtaining large-scale, physics-aware 3D assets, without requiring additional real-world data collection or computationally intensive generation. By combining the abundance of static 3D datasets with the strengths of vision-based perception models, our approach offers a scalable and generalizable pathway toward dynamic scene generation. The output of our pipeline can be seamlessly integrated into a wide range of renderers, simulators, and ecosystems, such as ROS, Isaac Sim, PyBullet, and Open3D, enabling flexible deployment across diverse downstream tasks. In summary, our key contributions are:
  • We present REACT3D, an efficient, automated workflow that leverages vision foundation models and vision-language models (VLMs) to recover object articulations from static 3D scenes and generate physically-enabled 3D scenes.
  • We provide a comprehensive evaluation of our pipeline against baseline methods.
  • We demonstrate the utility of our output assets across various platforms, including renderers, physics simulators, and the ROS ecosystem.
Figure 2: Overview of REACT3D. Given a static 3D scene, our method first applies open-vocabulary detection to identify openable objects and segmentation to extract their movable parts. We then estimate articulations and generate hidden geometry to obtain interactive objects. Finally, they are integrated with the static background to produce a simulation-ready interactive scene.
II RELATED WORK
II-A Articulation estimation
Articulated object motion estimation is a long-standing problem in computer vision and robotics [5, 6, 7]. Numerous works have explored predicting openable object parts and their motion parameters [6, 8, 9]. OPD [6] addresses this task by extending Mask R-CNN [10] to detect articulated parts and estimate their motions from single-view object images. OPDMulti [8] further generalizes OPD to handle real-world scenes containing multiple objects. In addition, 3DOI [9] proposes a transformer-based model to estimate the physical properties and affordance of objects using a single image and 2D query points as input.
Another line of work [7, 11, 12] focuses on predicting articulation directly on scene-level point clouds. SceneFun3D [11] introduces a large-scale dataset aimed at understanding object functionalities and affordances in real-world 3D environments, along with baseline models for predicting motion parameters associated with interactions involving functional elements (e.g., opening a drawer by pulling its handle). Articulate3D [12] builds upon ScanNet++ [2] annotated scenes to provide object-part hierarchy annotations and articulation parameters. However, scene-level point clouds often lack the fine-grained details of 2D images, and their annotations remain sparse. In this work, we propose an open-vocabulary approach to segment and articulate openable parts by leveraging off-the-shelf image-based articulation estimators together with segmentation foundation models.
II-B Articulated scenes from static observations
Early efforts on building interactive 3D scenes often required observing objects in multiple states or under user interaction. For instance, Ditto [13] reconstructs part geometry and joints by comparing 3D observations before and after human interaction. Similarly, PARIS [14] uses two sets of multi-view images captured under different articulation states to jointly recover part-level shape and motion parameters. Weng et al. [15] likewise leverage two RGB-D scans of an object at different articulations to learn a neural implicit twin with accurate kinematics. While these methods achieve faithful motion recovery, their reliance on paired observations or physical trials limits scalability. Some works have explored video-based static capture: e.g., ArtGS [16] employs 3D Gaussian splatting on casually captured video frames to reconstruct photorealistic objects with movable parts, but still benefits from careful segmentation or coarse templates to handle occluded regions. Overall, approaches that demand explicit state changes incur significant manual interaction overhead and scale poorly to large-scale scene conversion.
Recent research instead focuses on recovering articulated scenes directly from single-state observations using learned priors and foundation models. At the object level, Real2Code [17] proposes to generate programmatic code describing parts and joints given RGB images, using VLM to enforce structural consistency. Articulate-Anything [18] retrieves a similar 3D model from a shape library and infers plausible joints, enabling articulation for arbitrary objects but at the cost of fidelity to the actual geometry. Other methods tackle full scenes from minimal input: Digital Cousins [19] and URDFormer [20] generate an interactive scene from a single RGB image by leveraging prior knowledge. However, single-image approaches often yield inconsistent or non-faithful results, since a single view provides only partial evidence. More comprehensive pipelines integrate multi-view reconstruction and segmentation cues. DRAWER [4] reconstructs scenes from posed RGB frames using foundation models for part segmentation and articulation estimation, together with neural fields for appearance, achieving state-of-the-art realism and accurate articulations. These advances demonstrate a clear trend toward zero-shot articulated scene generation from static inputs, but challenges remain in balancing fidelity, physical accuracy, and scalability.
III METHOD
III-A Preliminaries
Problem formulation
Given a static 3D scene representation $\mathbf{S}$, our goal is to construct an interactive digital twin $\hat{\mathbf{S}}$ that augments the original geometry with articulated objects while preserving the static background (Fig. 2). The input $\mathbf{S}$ consists of a point cloud or a mesh $\mathbf{P}$ with per-vertex colors, captured by a registered RGB-D sensor array, together with a set of calibrated frames $\{(\mathbf{I}_{i},\mathbf{D}_{i},\mathbf{T}_{i},\mathbf{K}_{i})\}_{i=1}^{N}$, where $\mathbf{I}_{i}$ is an RGB image, $\mathbf{D}_{i}$ its depth map, $\mathbf{T}_{i}\in SE(3)$ the extrinsic pose, and $\mathbf{K}_{i}\in\mathbb{R}^{3\times 3}$ the intrinsic matrix. The output $\hat{\mathbf{S}}$ is decomposed as
$\hat{\mathbf{S}}=f(\mathbf{S})\;=\;\big(\,\mathbf{P}_{bg},\ \{\mathbf{O}_{j}\}_{j=1}^{M}\big),\quad\mathbf{O}_{j}=\big(\mathbf{p}_{j},\,\phi_{j},\,\mathbf{b}_{j}\big),$
where $\mathbf{P}_{bg}$ denotes the static background and each interactive object $\mathbf{O}_{j}$ is defined by its movable part $\mathbf{p}_{j}$, articulation parameters $\phi_{j}$, and completed hidden geometry $\mathbf{b}_{j}$. The articulation is parameterized as
$\phi_{j}=(t_{j},\mathbf{o}_{j},\mathbf{a}_{j},\rho_{j}),$
with joint type $t_{j}\in\{\mathrm{prismatic},\,\mathrm{revolute}\}$, origin $\mathbf{o}_{j}\in\mathbb{R}^{3}$, axis $\mathbf{a}_{j}\in\mathbb{R}^{3}$, and motion limits $\rho_{j}$.
Openable object
We define an openable object as one that contains a movable part actuated by either a prismatic or a revolute joint [21]. For a prismatic joint, the part translates along a fixed axis $\mathbf{a}$ with displacement $s\in[0,\rho]$, while for a revolute joint, it rotates about a unit axis $\mathbf{a}$ through origin $\mathbf{o}$ with angle $\theta\in[0,\rho]$. The motion of any point $\mathbf{x}$ belonging to the movable part in its closed configuration is
$\mathbf{x}^{\prime}\;=\;\begin{cases}\mathbf{x}+s\,\mathbf{a},&\text{prismatic},\\[6.0pt] R(\mathbf{a},\theta)(\mathbf{x}-\mathbf{o})+\mathbf{o},&\text{revolute},\end{cases}$
where $R(\mathbf{a},\theta)=\exp([\mathbf{a}]_{\times}\theta)$ denotes the axis-angle rotation and $\mathbf{a}]_{\times}$ is the skew-symmetric matrix of $\mathbf{a}$. This yields a compact representation of each movable part’s motion, parameterized by $(t,\mathbf{o},\mathbf{a},\rho)$, which serves as the basis for kinematic prediction.
III-B Openable object detection and segmentation
Interactive scene understanding requires semantic knowledge of which objects afford interaction (e.g., doors or drawers), beyond the appearance and geometry captured in static 3D scenes. As the first stage of our framework, openable object detection identifies candidate interactive instances, while part segmentation delineates their movable components from the surrounding structure to obtain aligned masks or meshes. This decomposes the input static scene into instance-level proposals that serve as the input domain for downstream articulation estimation, thereby determining the extent of interactivity in the resulting scene.
Open-vocabulary 2D detection and segmentation
Indoor scenes include diverse openable objects not captured by standard labels, motivating an open-vocabulary detection strategy to maximize interactive coverage. We therefore first process the RGB frames $\{\mathbf{I}_{i}\}_{i=1}^{N}$ of the static scene to recognize semantic tags of objects present in the scene using RAM++ [22]. Let $\mathcal{T}=\bigcup_{i=1}^{N}\texttt{RAM++}(\mathbf{I}_{i})$ denote the aggregated tag set. We then apply a VLM-based filtering method built on LLaVA [23] to retain only tags that satisfy our definition of openable objects, yielding $\mathcal{T}_{\mathrm{open}}=\{\,t\in\mathcal{T}\mid f_{\mathrm{VLM}}(t)=1\,\}$. Finally, Grounded SAM [24] is applied with prompts $\mathcal{T}_{\mathrm{open}}$ on each frame $\mathbf{I}_{i}$ to localize and segment movable parts, producing a set of 2D instance masks $\{\mathbf{m}_{i,k}\}_{k=1}^{n_{i}}$, where $n_{i}$ denotes the number of detected movable parts in frame $\mathbf{I}_{i}$, which are post-processed with hole filling to close interior gaps, as Grounded SAM can miss door handles. Unlike prior approaches restricted to predefined labels, our open-vocabulary method mitigates label bias and yields better coverage of long-tail openable objects.
2D-to-3D segmentation via multi-view fusion
Zero-shot 3D instance segmentation methods [25] operating on point clouds or meshes rarely yield reliable part-level segmentation. We therefore adopt a 2D-to-3D multi-view fusion paradigm, similar to DRAWER, to lift robust 2D masks into 3D. For each frame, we construct the model-view-projection matrix from $\mathbf{T}_{i},\mathbf{K}_{i}$ to recover the camera view. The scene mesh is then rasterized under each view, and the front-most faces covered by each mask are selected as its 3D projection. To recover the full structure of each object, as they are not always fully visible from a single view, we fuse the per-mask projections by building a face graph and applying Louvain community detection [26]. Unlike DRAWER, which retains only one mask above the Intersection over Union (IoU) threshold for each fusion result, our approach ranks the masks by IoU and preserves the top-$k$ views that exceed the threshold, providing multiple perspectives that improve the robustness of subsequent articulation estimation. In addition, we replace its VLM-based filtering of part masks, which discards candidates without visible handles, with an alternative validation step at the articulation estimation stage (Sec. III-C). This design allows the detection of openable objects even when handles are absent or visually nonsalient, as shown by the comparison in Fig. 6. We then re-segment each object by invoking SAM with point prompts seeded from previously fused 3D proposals.
Another challenge in 2D-to-3D lifting is that segmentation often produces the whole object rather than its movable part, e.g., the whole cabinet instead of just the door. To address this, we refine each segmented 3D mesh by applying RANSAC [27] to estimate an approximately vertical, finite-thickness plane and retain the estimate with the largest contour area. Naive planar clipping, however, may fail to capture protruding handles. To preserve such structures while discarding static body regions, let $\mathbf{n}$ denote the normal of the estimated plane and $\mathbf{v}_{i}$ the viewing direction derived from the camera pose $\mathbf{T}_{i}$ of the frame that produced the segmentation. We choose the inward-pointing normal such that $\langle\mathbf{n},\mathbf{v}_{i}\rangle>0$, and discard the geometry behind the plane along this direction. This produces high-quality 3D movable-part meshes $M^{\mathrm{obj}}_{j}=(V_{j},F_{j})$, where $V_{j}$ and $F_{j}$ denote the vertex and face sets of the segmented mesh, together with the top-$k$ RGB views and their associated masks $\{\mathbf{m}_{j,i}\}_{i=1}^{k}$. The resulting part meshes are spatially aligned with the input static scene and provide the foundation for generating interactive objects.
III-C Articulation estimation
To enable meaningful manipulation, each movable part must be endowed with kinematic parameters that define its motion type and limits. Accordingly, for every movable part, we estimate an articulation tuple $\phi(m)=\big(t,\mathbf{o},\mathbf{a},\rho\big)$. In our framework, we leverage OPDMulti [6, 8] to obtain initial articulation estimates from a single image, producing masks of movable parts together with their joint type, origin, axis, and motion range.
Candidate filtering
Since OPDMulti can detect multiple movable parts from a single RGB frame, it offers the opportunity to further validate the candidate openable objects obtained from the segmentation stage. For each segmented 3D mesh, we utilize the pre-saved top-$k$ best-view images and their SAM-generated masks $M_{\mathrm{SAM}}=\{\mathbf{m}_{i}\}_{i=1}^{k}$. We then run OPDMulti on these views and compare its predicted movable-part masks $M_{\mathrm{OPDM}}$ against $M_{\mathrm{SAM}}$. A candidate is considered valid if the IoU between the two masks exceeds a threshold in at least one of the top-$k$ views; otherwise, the 3D mesh is discarded as an invalid case.
Articulation refinement
While OPDMulti provides a reasonable initial articulation, direct application to interactive objects may fail to reproduce realistic joint behavior. To recover faithful interactions, we refine the initial estimates for each validated movable part by computing an oriented bounding box (OBB) of the part mesh with COMPAS [28]. The OBB is denoted $(\mathbf{U},\mathbf{c},\mathbf{s})$, where $\mathbf{U}=[\mathbf{u}_{1}\,\mathbf{u}_{2}\,\mathbf{u}_{3}]$ is a set of orthogonal edges, $\mathbf{c}$ is the center, and $\mathbf{s}=(s_{1},s_{2},s_{3})$ are the side lengths with $s_{1}\geq s_{2}\geq s_{3}$. Since openable parts are typically plate-like, we define the front face $P_{\mathrm{front}}$ as the OBB face spanned by the two longest edges $\mathbf{u}_{1}$ and $\mathbf{u}_{2}$ with unit normal $\mathbf{n}_{\mathrm{front}}=\mathbf{u}_{1}\times\mathbf{u}_{2}$. Let the initial articulation be $\phi(m)=(t,\mathbf{o},\mathbf{a},\rho)$. For prismatic joints, we update only the joint axis direction by aligning it with the front-face normal,
$\mathbf{a}^{\prime}\;=\;\operatorname{sign}\!\big(\langle\mathbf{a},\mathbf{n}_{\mathrm{front}}\rangle\big)\,\mathbf{n}_{\mathrm{front}},\qquad\phi^{\prime}(m)=(t,\mathbf{o},\mathbf{a}^{\prime},\rho),$
which preserves the origin $\mathbf{o}$ while enforcing translation orthogonal to $P_{\mathrm{front}}$. For revolute joints, we first select the in-plane principal direction most consistent with the initial axis,
$\mathbf{l}_{s}\;=\;\arg\max_{\mathbf{v}\in\{\mathbf{u}_{1},\mathbf{u}_{2}\}}\,\big|\langle\mathbf{a},\mathbf{v}\rangle\big|,\qquad\mathbf{a}^{\prime}\;=\;\operatorname{sign}\!\big(\langle\mathbf{a},\mathbf{l}_{s}\rangle\big)\,\mathbf{l}_{s}.$
Next, we place the origin by comparing the joint axis line defined by the initial axis $L(\mathbf{o},\mathbf{a})=\{\mathbf{o}+s\,\mathbf{a}\mid s\in\mathbb{R}\}$ to the two OBB front-face edges that are parallel to $\mathbf{l}_{s}$. We compute the shortest distance from $L$ to each of these edges, select the closer one, and take its midpoint. This point is then projected along $\mathbf{n}_{\mathrm{front}}$ onto the mid-surface equidistant between $P_{\mathrm{front}}$ and its parallel opposite OBB face $P_{\mathrm{back}}$, where the projected point serves as the refined joint origin $\mathbf{o}^{\prime}$. This procedure anchors the origin to the mid-surface edge most consistent with the initial axis while ensuring the most appropriate location within the part’s thickness. Finally, we update
$\phi^{\prime}(m)=(t,\mathbf{o}^{\prime},\mathbf{a}^{\prime},\rho).$
Intuitively, the prismatic axis is made orthogonal to the part face, whereas the revolute axis is aligned with the dominant in-plane edge and centered within the part thickness; in both cases the axis sign is chosen to minimize angular deviation from the initialization. This refinement aligns the estimated kinematic parameters with the observed geometry of the part, yielding more realistic motion.
III-D Hidden geometry generation
The input static 3D scene encodes only the visible surface geometry of objects. Consequently, even though articulation estimation enables object interaction, the interior behind the movable part remains unmodeled upon opening. This creates visually unrealistic artifacts and, more critically, limits practical downstream applications such as robotics, where the interior volume is essential for manipulation. To construct a complete interactive object, we generate the hidden geometry through cavity completion, producing consistent box-like structures that approximate the container behind each movable part. Fig. 3 demonstrates the process of interactive object generation.
As in Sec. III-C, we compute an OBB for each movable part and take the face formed by its two longest edges as the front face $P_{\mathrm{front}}$. These edges define the width and height of the inner box, while its depth is measured along the inward-pointing normal of the front face, $\mathbf{n}_{\mathrm{front}}=\mathbf{u}_{1}\times\mathbf{u}_{2}$. We define the inner-box depth as
$d_{\mathrm{in}}\;=\;\min\big(\,d_{\mathrm{image}},\;d_{\mathrm{hit}},\;d_{\mathrm{mesh}}\,\big),$
and then instantiate a rectangular cavity whose front coincides with $P_{\mathrm{front}}$ and whose extent is $d_{\mathrm{in}}$ along $\mathbf{n}_{\mathrm{front}}$. Here, $d_{\mathrm{image}}$ is an image-derived bound given by the farthest background depth in the RGB-D frame used to segment the part mesh, projected onto $\mathbf{n}_{\mathrm{front}}$. $d_{\mathrm{hit}}$ is obtained by probing along $\mathbf{n}_{\mathrm{front}}$ from the OBB centroid to the nearest scene intersection, and taking its depth if a supporting plane can be fitted there using RANSAC. Finally, $d_{\mathrm{mesh}}$ is the boundary depth of the static scene mesh, serving as an upper bound to prevent leakage outside the scene.
Predicted 2D Mask Segmented Mesh Predicted Joint Refined Joint Interactive Object Opened Object
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Pipeline for interactive object generation. From left to right, the figure shows key intermediate results of interactive object generation. In the last column, the thin red line highlights the contour of the base part.
III-E Interactive Scene Integration
Our goal is to produce a complete interactive scene that can be directly used in downstream applications. The output includes interactive objects consisting of movable parts with refined articulation and completed internal geometry, all without duplicates, together with the textured static background integrated into a unified representation.
Duplicate removal
To ensure clean assembly of the final interactive scene, we perform a last deduplication pass. Let $\{S_{i}\}$ denote movable-part meshes with point sets $\{X_{i}\}$. We compute the IoU directly on these point sets, $\mathrm{IoU}(X_{i},X_{j})=\tfrac{|X_{i}\cap X_{j}|}{|X_{i}\cup X_{j}|}$. The procedure has two stages. Stage 1 (pairwise pruning): for every pair $(i,j)$, if $\mathrm{IoU}(X_{i},X_{j})\geq\tau_{\mathrm{dup}}$, we keep the mesh with greater information content (measured by point count) and discard the other, producing a reduced set $\{\hat{S}_{i}\}$ with point sets $\{\hat{X}_{i}\}$. Stage 2 (subdivision pruning): duplicates can also appear when a larger mesh can be represented by the union of several smaller, complementary meshes while no single pair exceeds $\tau_{\mathrm{dup}}$. We sort $\{\hat{S}_{i}\}$ by $|\hat{X}_{i}|$ in descending order and iterate from large to small so that larger parts can be explained by smaller ones first. For each $\hat{S}_{i}$, we collect candidate smaller meshes $\mathcal{C}_{i}=\{\,j:|\hat{X}_{j}|<|\hat{X}_{i}|,\ \mathrm{IoU}(\hat{X}_{i},\hat{X}_{j})\geq\tau_{\mathrm{low}}\,\}$, where $\tau_{\mathrm{low}}$ is a small IoU threshold, and enumerate their combinations. If there exists a subset $\mathcal{U}\subseteq\mathcal{C}_{i}$ such that $\mathrm{IoU}\!\left(\hat{X}_{i},\ \bigcup_{j\in\mathcal{U}}\hat{X}_{j}\right)\geq\tau_{\mathrm{dup}}$, we remove $\hat{S}_{i}$ as a redundant superset. In the end, we remove all redundant movable parts together with their associated inner boxes before assembling the final interactive scene.
Scene assembly
We integrate interactive objects into the original static scene by carving the background geometry using the world-frame OBBs of deduplicated movable parts. Let the original static scene mesh be $\mathbf{P}$ and the set of movable-part OBBs be $\{B_{j}\}$. To remove residual fragments from imperfect segmentation and spurious points in the original scan, we delete all vertices lying inside any $B_{j}$ together with their incident faces. This yields a cleaned background mesh $\mathbf{P}_{\mathrm{rem}}\;=\;\mathbf{P}\setminus\bigcup_{j}B_{j}$. We then assemble the final interactive scene by combining $\mathbf{P}_{\mathrm{rem}}$ with the interactive objects, ensuring a clean integration without interpenetration.
Texture generation
As both the original scene mesh and the segmented movable-part meshes provide only per-vertex colors, we generate texture maps based on InstantTexture [29]. We first perform UV unwrapping with Xatlas [30] on the movable parts and the cleaned remainder mesh. Given the UV parameterization, we rasterize each triangle in atlas space and compute per-texel colors by barycentric interpolation of the source vertex colors, yielding dense textures aligned to the unwrapped charts. We then repair uncovered texels caused by sampling gaps or occlusions via inpainting, followed by smoothing with low-pass filtering to attenuate aliasing across chart boundaries. Finally, the generated textures are bound to their corresponding meshes and, together with the refined articulation parameters, exported as .dae assets that are referenced in the scene description files.
Export and simulator support
To ensure portability across visualization and simulation stacks, we export each interactive scene in two widely adopted formats: URDF for the robotics ecosystem and USD for graphics/physics engines. The exports package the cleaned background, articulated interactive objects, and textures, enabling direct deployment in PyBullet [31], ROS [32], and Isaac Sim [33]. Fig. 4 shows the simulated interactive scenes generated by REACT3D from different inputs in Isaac Sim. For ROS, we additionally provide a lightweight GUI for manipulating the interactive scene during visualization and benchmarking (see Fig. 5).
TABLE I: Quantitative results on openable object detection. Openable object detection on 30 ScanNet++ scenes rich in openable objects. We report Precision, Recall, and F1 at IoU thresholds $\tau\!\in\!\{0.25,0.50\}$. For URDFormer, $\tau$ denotes the coverage ratio, while for other methods it denotes IoU. For URDFormer, the number reflects only openable objects visible in the selected region with the highest density of openable objects, since the method takes a single RGB image as input.
Method Representation #Openable Objects $\tau^{*}=0.25$ $\tau^{*}=0.50$
Precision $\uparrow$ Recall $\uparrow$ F1 Score $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ F1 Score $\uparrow$
URDFormer Mesh 226 $0.429$ $0.226$ $0.296$ $0.361$ $0.190$ $0.249$
DRAWER (GT-mesh) Points & Mesh 540 $0.434$ $0.141$ $0.213$ $0.326$ $0.106$ $0.159$
Ours Points & Mesh 540 0.731 0.393 0.511 0.666 0.357 0.465
Static Scene
REACT3D
Figure 4: Qualitative results of REACT3D. Static input scenes from ScanNet++ and the interactive outputs generated by REACT3D, visualized in Isaac Sim.
ROS
Isaac Sim
Figure 5: Manipulation GUIs. Interfaces in ROS and Isaac Sim enabling per-object articulation control and benchmarking.
IV Experiments
IV-A Experimental Setup
Datasets
To comprehensively evaluate our method and baselines, we use ScanNet++ [2], a widely adopted dataset that provides high-quality posed RGB-D reconstructions across diverse indoor environments. We select 30 scenes rich in objects with potential articulation, ranging from single rooms to full apartments. For quantitative evaluation, we adopt Articulate3D [12] as ground truth, which augments ScanNet++ with part-level semantic labels and articulation annotations (joint type, origin, axis, and motion range). Our evaluation focuses on objects with open-close motion, while small controls such as buttons and knobs are excluded. For benchmarking, we evaluate our method against the baselines URDFormer and DRAWER.
Evaluation metrics
We assess performance along two dimensions: movable-part detection and articulation parameters. For detection, we compute vertex-level IoU over all pairs of predicted movable-part meshes and ground-truth part meshes from Articulate3D in world coordinates. As the predicted parts are segmented directly from the input static scene, both predictions and ground truth reside in the same coordinate frame, requiring no additional alignment. A prediction is counted as a true positive if its IoU with at least one ground-truth part exceeds $\tau$. We report precision, recall, and F1 at $\tau\in\{0.25,0.50\}$ across all scene objects.
For articulation, we consider only true positive detections, evaluating each predicted joint against its matched ground-truth counterpart. We first assess the predicted joint type and report joint-type accuracy. Following the MultiScan protocol [7], articulation is evaluated for parts with a correct type match using two metrics: (i) Minimum Distance (MD), the shortest distance between the predicted and ground-truth joint lines, which better captures positional error than origin-to-origin distance since origin differences along the joint axis do not alter the induced motion; and (ii) Orientation Error (OE), the angle between the predicted and ground-truth joint axes, capturing directional error. As prismatic motion is origin-invariant, we report only OE for prismatic joints, while both MD and OE are reported for revolute joints.
To ensure fair comparison across methods with different articulation priors, we adopt a three-way joint-type taxonomy: prismatic, horizontal revolute, and vertical revolute. Most indoor revolute mechanisms are near-vertical, so methods that always predict vertical axes gain an artificial advantage if the revolute joint is treated as a single class in evaluation, yielding deceptively low OE. Conversely, methods that explicitly model horizontal axes can be penalized with errors near $90\text{\,}\mathrm{\SIUnitSymbolDegree}$. To mitigate this bias, we classify a revolute joint as horizontal if its axis forms less than $45\text{\,}\mathrm{\SIUnitSymbolDegree}$ with the ground plane, and vertical otherwise. Joint-type accuracy and OE are reported under this taxonomy, revealing true motion behavior while isolating differences in method capability.
Input
DRAWER (GT-mesh)
REACT3D
Figure 6: Qualitative comparisons for scene-level movable-part detection and articulation estimation. Red arrows indicate revolute joints, while blue arrows indicate prismatic joints. In (a), our method achieves significantly higher performance than DRAWER when openable objects lack visible handles. In (b), DRAWER misclassifies drawers as revolute joints, whereas our method provides correct predictions. Visualizations are generated with Open3D [34].
Input
URDFormer
DRAWER
REACT3D
Figure 7: Qualitative comparison of generated interactive scenes. For each input static scene, we show results produced by each native method without any adaptation or modification. Visualizations follow each method’s native support: URDFormer is rendered in PyBullet, while DRAWER and REACT3D are rendered in Isaac Sim. Input corresponds to the ScanNet++ scene mesh, the posed RGB-D frames, or a captured image of the scene.
TABLE II: Quantitative results on articulation estimation and MOD evaluation. Articulation evaluation on the same 30 scenes. Joint type accuracy is computed over correctly detected movable parts, while minimum distance (MD) and orientation error (OE) are reported only for parts with correctly predicted joint types. Conditions are the same as in Tab. I.
Method
#Openable
Objects
$\tau^{*}=0.25$ $\tau^{*}=0.50$
[…]
URDFormer 226 $50.9$ $0.583$ 0.500 $22.6$ $11.5$ $11.5$ $1.3$ $46.5$ $0.491$ 0.609 $19.0$ $8.9$ $8.9$ $1.8$
DRAWER (GT-mesh) 540 $73.7$ $0.288$ $1.621$ $14.1$ $10.4$ $10.0$ $5.2$ $78.9$ $0.313$ $1.714$ $10.6$ $8.3$ $8.0$ $3.7$
Ours 540 91.5 0.203 $1.161$ 39.3 35.9 35.9 24.8 92.2 0.180 $1.130$ 35.7 33.0 33.0 23.5
IV-B Baselines
URDFormer
URDFormer generates an interactive scene from a single RGB image without guaranteed metric scale or an explicit mapping to the source 3D scene. For fair comparison, we adopt a concise placement-scaling protocol. For each static scene, we select a region rich in potentially interactive objects and capture a posed RGB-D image. We run the official model, extract per-object 3D bounding boxes and a global box, estimate a global scale $s$ as the ratio between the vertical extent of the ground-truth region and that of the predicted global box, then scale and place the prediction back into the static scene using the pose and depth.
As the output is not segmented from the scene mesh, vertex-level IoU is inapplicable. We therefore evaluate detection via 3D coverage: for a predicted box $B$ and a ground-truth movable-part mesh $M$ with vertices $V(M)$ in world coordinates, $\mathrm{cov}(B,M)=\frac{|V(M)\cap B|}{|V(M)|}$ and a part is detected if $\max_{B}\mathrm{cov}(B,M)\geq\tau$. Articulation is then evaluated with the same protocol and metrics defined above.
DRAWER
DRAWER reconstructs a scene mesh from RGB frames via monocular depth and normal estimation and generates an interactive scene on top. For fair comparison on ScanNet++, we bypass reconstruction and directly use ground-truth meshes as input. All downstream modules follow official settings and pretrained weights, except that the original GPT-4o component is replaced with the same LLaVA model used in our framework. We denote this configuration as DRAWER (GT-mesh). This protocol aligns the input domain with ours, ensuring independence from reconstruction errors in evaluation.
IV-C Results
Openable object detection
We first assess how many openable objects each method recovers in the generated interactive scenes. Table I compares URDFormer, DRAWER (GT-mesh), and REACT3D under the experimental setup. Note that the output scenes of URDFormer are not at true scale as explained in Sec. IV-B, whereas the other methods produce scenes at true scale. Across all metrics, our framework achieves the best results and substantially outperforms both URDFormer and DRAWER, demonstrating its superior ability to consistently and accurately recover interactive objects across diverse indoor scenes.
Articulation estimation
As shown in Tab. II, our approach achieves over 90% joint-type accuracy, substantially surpassing all baselines. It also attains the lowest MD and OE among methods with learned articulation, outperforming DRAWER, the prior state of the art with true-scale results. Although URDFormer reports the lowest OE, this is due to its heuristic strategy of presetting articulation axes to vertical or horizontal relative to the ground based on joint type. Fig. 6 shows scene-level comparisons of movable-part detection and articulation results, while Fig. 7 compares interactive scenes generated by the native methods.
We further report results under the MOD setting in Tab. II, following the same experimental setup. This setting provides a cumulative view of articulation parameter correctness and reveals which metric constitutes the primary performance bottleneck. Here, PDet measures detection accuracy, while +M, +MO, and +MOD denote progressively stricter criteria: correct motion type, motion type with OE $10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, and additionally MD $0.25\text{\,}\mathrm{m}$\$, respectively [7]. Each stage is evaluated only if all preceding conditions are satisfied.
Ablation study
To assess the contribution of our articulation refinement module, we conduct an ablation study by comparing results with and without it. As shown in Tab. III, refinement consistently improves performance across both IoU thresholds. The reduction in minimum distance demonstrates more accurate joint localization, while the dramatic decrease in orientation error underscores its importance for reliable axis direction estimation. Overall, the results demonstrate that refinement is crucial for producing precise articulation parameters and achieving robust interactive scene generation.
TABLE III: Ablation study on articulation refinement. We report minimum distance (MD) and orientation error (OE) at IoU thresholds $\tau\!\in\!\{0.25,0.50\}$. Refinement consistently enhances accuracy, with notably large improvements in OE.
Method $\tau=0.25$ $\tau=0.50$
[…]
w/o refinement $0.226$ $6.721$ $0.213$ $6.638$
w/ refinement 0.203 1.160 0.180 1.130
V Conclusion
We presented REACT3D, a scalable zero-shot framework that transforms static 3D scenes into interactive digital twins with articulated objects. By combining open-vocabulary detection, articulation refinement, hidden-geometry completion, and seamless simulator integration, REACT3D achieves state-of-the-art performance on scene-level openable-object detection and articulation estimation. Our results demonstrate the practical utility of REACT3D in generating simulation-ready assets, providing a foundation for large-scale research in articulated scene understanding and embodied intelligence.